ramon is 3 years older than sally. four times ramon's age ten years ago is 5 less than twice ramon's age now. how old are they now?
Answer by MathTherapy(10552)

Let Sally’s age be s
Then Ramon’s age now is s + 3
Ten years ago, Ramon was s + 3 – 10 = s – 7
If ten years ago, 4 times Ramon’s age was 5 less than twice his age now, then, 4(s – 7) = 2(s + 3) - 5
4s - 28 = 2s + 6 - 5
2s = 29
s = 14.5

Sally is now 14.5, which means that Ramon is now 17.5 (14.5 + 3).

Checking:
4(17.5 – 10) = 2(17.5) – 5
4(7.5) = 35 - 5
30 = 30